# Identifying User Goals from UI Trajectories

### v2

@@ 5.1 Human Performance Evaluation
To asses task difficulty and establish a baseline for models, two NLP practitioners, unfamiliar with the datasets, independently composed 50 task descriptions from Mind2Web, and 50 from AitW. An annotator then evaluated [^1] During the evaluation it was observed that human-generated tasks were inherently fulfilled by the trajectory. Therefore, we do not report this score. these descriptions as explained in § 3.2.
@@
The results, summarized in Table 1, reveal more matches between the human annotators and the gold task descriptions in Mind2Web compared to AitW. Upon analysis of the AitW disagreements, we found that the gold descriptions were often more specific than those provided by human annotators. In some cases, the gold tasks were already fulfilled by the trajectory, resulting in no clear interactions that indicate the user goal. For example, if the task was “Turn WiFi on” and the WiFi was already on, the annotators inferred a more general task, such as “Show WiFi settings”. Further analysis of non-matching records across both datasets revealed that ambiguous task descriptions contributed to disagreements between raters. Detailed analysis and examples can be found in Appendix C.
5.2 Model Performance Evaluation
Model evaluation included manual assessment of 100 predicted task descriptions and automated evaluation of 1,000 predictions per model and dataset (details provided in Appendix B).
@@ 5.2 Model Performance Evaluation
The results, outlined in Table 2, reveal that Gemini outperforms GPT-4 on both datasets, achieving higher match scores but still falling short of human performance. In AitW, GPT-4 often generated too abstract goals such as “Explore the main page of a shopping app” resulting in higher non-match scores. In contrast, Gemini tended to be more specific and heavily influenced by the content on the screen. Both models, in certain instances, misidentified the actual user intent (e.g. “Watch top rated movie trailer” vs. “Watch The Dark Knight trailer”) and exhibit limitations in visual screen understanding, leading to missing details, incorporating irrelevant information and hallucinations of non-existent information. We provide a detailed error analysis and further discussion in Appendix C.
Experiments with few-shot prompting did not yield improved results. We hypothesize that the large context size, resulting from including multiple trajectories with several images each in the prompt, were too complex models to leverage.
6 Conclusion and Future Work
@@
In this paper, we introduced a novel task for UI-related agents by identifying user goals from UI trajectories and proposed a reliable evaluation methodology. Our experiments on Android and web datasets reveal a significant gap between humans and state-of-the-art models, with a detailed analysis provided for this discrepancy.
@@
Future work may include fine-tuning experiments and enhancing visual understanding for UI environments. We propose testing our task by evaluating how well it improves downstream tasks like agent personalization and completions. Additionally, expanding the scope to other GUIs like iOS and Windows would broaden the applicability and impact of our findings.
@@ Appendix A Datasets Overview
Figure 2: An instance from Mind2Web, representing a full trajectory accomplishing the task description above.
Figure 3: An instance from AitW, representing a full trajectory accomplishing the task "Set an alarm". The blue plus sign indicates the area on the screen where the tap occured.
Mind2Web
@@
Android-In-The-Wild
The Android in the Wild (AitW) dataset stands out due to its extensive variety of tasks, covering 4 domains: Google-Apps, Install, Web-Shopping, and General, as well as a single-step domain excluded from this paper’s analysis. The dataset consists of a substantial collection of high-level instructions, trajectories of varying lengths, and a notable variety of apps and websites. A key strength of the AitW dataset lies in the diversity of task instructions, which provided in-depth insights into the complexities of understanding user goals across different domains.
The high-level goal instructions were sourced from various sources, including humans (both crowd-sourced and the authors themselves), LLM-generated prompts, and technical documentation such as PixelHelp Li et al. (2020). The creation process involved human annotators performing tasks on Android emulators, with their gestures being recorded.
The AitW dataset contains 700k+ episodes comprising 30k unique instructions. These episodes were recorded on mobile devices running four different versions of Android. Each episode contains natural language instructions and observation-action pairs. The observations are screenshots, while the actions are one of three types: tap, drag, or typing. Gesture actions are represented as taps and drags at specific <x,y> coordinates on the screen.
@@ Appendix B Experiments
Data
For manual evaluation, we sampled 100 random data points from the test sets of both AitW and Mind2Web. For automatic evaluation, we used the entire Mind2Web test set, consisting of 1013 examples, and randomly sampled 250 examples from each of the four domains in AitW, resulting in a total of 1000 examples (as described in Appendix A).
Web
@@
To address this, we implemented a heuristic truncation method, reducing image height while ensuring the interaction element remained visible within the truncated image. This was achieved by utilizing the bounding box metadata provided by Mind2Web. Similar to the approach taken by Zheng et al. (2024) and introduced by Yang et al. (2023a), we drew the red bounding box around the interaction element to guide the model’s attention. Additionally, adding the bounding box helped resolve ambiguities at times that the textual action description is not sufficient. For example, sometimes action descriptions are simply empty, and does not contain any description about the element itself. While in other cases, action descriptions exist, but the textual description matches multiple element descriptions and thus results in ambiguity that is only resolved by drawing the bounding box. For example, a button labeled "Add to Cart" is typically associated with each item in a web shopping list. Without the bounding box, it is impossible to determine which specific button was clicked.
Android
@@
To address these challenges, we utilized the dataset’s utilities to overlay actions on top of the screenshots, as well as bounding box annotations for post-process detected UI elements. However, we found that the added element annotation marks often confused the model added noise and critical information on the screen. As a result, we proceeded with experiments using a sequence of screenshots that had the actions (tap, drag, and type) drawn over them. AitW’s provided visualization tools to draw the actions, also included labels of special actions such as the back button, home button, and enter. As well as a special "status" action: either Task Complete or Task Impossible. However, we found that the label Task Complete often confused the models. Despite our efforts to make the models ignore it, we eventually abandoned this specific annotation as it is a technicality of data representation.
@@ Appendix C Error Analysis
Figure 4: An illustration in which the user chose a specific car primarily for its 12-inch feature, but since it was also the cheapest, annotators incorrectly assumed cost was the deciding factor.
@@
Figure 5: Comparison of "Match" proportions between Gemini 1.5 Pro and GPT-4-Turbo models across different AitW domains
Android
Comparing GPT and Gemini’s performance on the Android dataset revealed notable variations across different domains. Gemini consistently surpassed GPT in all domains, primarily due to GPT’s tendency to offer abstract goals and frequent random hallucinations. Gemini’s errors varied based on the specific domain.
Both models were proficient in the “General” domain, but faced challenges in “Web Shopping” and “Google Apps”. This disparity is due to the nature of the “General” dataset, which contains search queries that explicitly reveal the user’s intent.
“Google Apps” tasks overlap with PixelHelp, primarily involving settings configuration. Some tasks are ambiguous, such as a button toggle, but the trajectory displays only one option. Additionally, specific tasks request actions that have already been completed, e.g Turn On location history, but the trajectory only shows viewing the Location History setting page, leading to confusion for models. Additionally, the models faced difficulty comprehending the correct order of the sequence of actions that occurred, frequently mistaking the final state (on or off).
@@
The “Install” domain often presents ambiguous tasks in the format “open (install if not installed)” which confuses both models. Furthermore, in some cases the apps were already pre-installed which made it impossible to predict (model predicts “open”), providing only partial satisfaction in one direction.
These results indicate that further refinement and training may be needed to improve the models’ performance in specific domains.
@@ Appendix D Instructions and Prompts
Figure 6: Instructions for human-annotators to conclude if a task is fulfilled by a trajectory, and if two task descriptions satisfying each other. These instructions were the core prompt of the automatic evaluator.
Figure 7: Goal identification annotation task instruction to measure human performance.
Figure 8: Instruction used to guide GPT and Gemini when predicting a task description given a UI trajectory. We swap AitW section and Mind2Web according to the input.
Figure 9: Model instructions for evaluating Satisfaction relation between two task descriptions (A and B), given a corresponding trajectory. Includes web/mobile data format instructions and few-shot examples for experimentation (detailed on the next page).
Figure 10: Few-Shot exemplars demonstrating the Satisfaction Relation Prompt (Figure 9). These simplified examples highlight the nuances of task satisfaction in real-world scenarios.

### v3

@@ 6.1. Human Performance Evaluation
To asses task difficulty and establish a baseline for models, two NLP practitioners, unfamiliar with the datasets, independently composed 50 task descriptions from Mind2Web, and 50 from AitZ. An annotator then evaluated [^1] During the evaluation it was observed that human tasks were inherently fulfilled by the trajectory. them as explained in § 4.2.
@@
The results, summarized in Table 1, reveal more matches between the human annotators and the gold task descriptions in Mind2Web compared to AitZ. Upon analysis of the disagreements, we found that the gold descriptions in Android were often more specific than those provided by human annotators. In most cases, the gold tasks were already fulfilled by the trajectory, resulting in no clear interactions that indicate the user goal. For example, if the task was “Turn WiFi on” and the WiFi was already on, the annotators inferred a more general task, such as “Show WiFi settings”. Further analysis of non-matching records across both datasets revealed that ambiguous trajectories caused human disagreements. Detailed analysis in Appendix C.
6.2. Model Performance Evaluation
Model evaluation included manual assessment of 100 predictions from each dataset and automatic evaluation using the entire test sets: 1,013 from Mind2Web and 506 from AitZ (see Appendix B).
@@ 6.2. Model Performance Evaluation
The results, outlined in Table 2, reveal that Gemini outperforms GPT on the Mind2Web dataset, achieving higher match scores but still falling short of humans. In AitZ, both models perform comparably. Gemini tends to be more specific and detailed than the gold references, while GPT often generates more general, abstract goals. Both models, in certain instances, misidentified the actual user intent (e.g. “Watch top rated movie trailer” vs. “Watch The Dark Knight trailer”) and exhibit limitations in visual screen understanding, leading to missing details, incorporating irrelevant information and hallucinations of non-existent information. Detailed discussion in Appendix C.
These results underscore the complexity of accurately capturing user intent. Our experiments also concluded few-shot learning. However, in our setting, this is challenging due to multiple images, textual inputs and the thought process within example. Adding a single example deteriorated results, likely because the large context size and the difference between the exemplar and the test trajectory.
7. Conclusion and Future Work
@@
In this paper, we introduced the novel task of identifying user goals from UI trajectories and proposed a reliable evaluation methodology. Our experiments on Android and web datasets reveal a significant gap between humans and state-of-the-art multi-modal models, which is analyzed in detail.
@@
Future work may include fine-tuning experiments and enhancing visual understanding for UI environments. We propose testing our task by evaluating its utility on downstream tasks like agent personalization and suggestions. Additionally, expanding the scope to other GUIs like iOS and Windows would broaden the impact of our findings.
@@ Appendix A Datasets Overview
Figure 2. An instance from Mind2Web, representing a full trajectory accomplishing the task description above.
[…]
The figure depicts a sequence of user interactions within a web environment to complete the task of ”Find unique experiences in London of maximum one hour duration which are rated four stars and above”. It consists of six sub-images, each showcasing an web screen representing a specific step in the user’s actions. This example is taken from the Mind2Web dataset.
Figure 3. An instance from AitW, representing a full trajectory accomplishing the task ”Set an alarm”. The blue plus sign indicates the area on the screen where the tap occurred.
[…]
The figure depicts a sequence of user interactions within an Android environment to complete the task of ”Set an Alarm.” It consists of five sub-images, each showcasing an Android screen representing a specific step in the user’s actions. This example is taken from the Android In the Wild dataset.
A.1. Web
@@
A.2. Android
The Android in the Wild (AitW) dataset stands out due to its extensive variety of tasks, covering 4 domains: Google-Apps, Install, Web-Shopping, and General, as well as a single-step domain excluded from this paper’s analysis. The dataset consists of a substantial collection of high-level instructions, trajectories of varying lengths, and a notable variety of apps and websites. The Android in the Zoo (AitZ) dataset was sampled from AitW to reduce tasks redundancy and to filter erroneous data points, resulting in more unique and higher-quality task descriptions.
The high-level goal instructions in AitW were sourced from various sources, including humans (both crowd-sourced and the authors themselves), LLM-generated prompts, and technical documentation such as PixelHelp (Li et al., 2020). The creation process involved human annotators performing tasks on Android emulators, with their gestures being recorded.
The episodes were recorded on mobile devices running four different versions of Android. Each episode contains natural language instructions and observation-action pairs. The observations are screenshots, while the actions are one of three types: tap, drag, or typing. Gesture actions are represented as taps and drags at specific ¡x,y¿ coordinates on the screen.
@@ B.1. Models Configuration
B.2. Data
For the manual evaluation, we randomly sampled over 100 data points from the both Android and Web datasets. We then prompted models to predict the user goal. During the evaluation, we conducted a thorough verification process to ensure that both the gold references and the trajectories were of high quality.
We rejected instances where the trajectory did not fulfill the gold reference, which occurred more frequently for Android. Additionally, for web-based tasks, we excluded cases where the screenshot was not rendered properly, making it difficult to interpret the user’s action, as well as instances where the bounding box was empty.
As a result of this verification process, 17% of the Android examples and about 5% of the web examples were rejected. Ultimately, as a result of this process, the samples that were manually evaluated in Sections 4.2 and 6 were of high quality and free from such issues. For the automatic evaluation, as described in 6, we utilized the entire test sets from both datasets.
B.3. Web
@@
To address this, we implemented a heuristic truncation method, reducing image height while ensuring the interaction element remained visible within the truncated image. This was achieved by utilizing the bounding box metadata provided by Mind2Web. Similar to the approach taken by (Zheng et al., 2024) and introduced by (Yang et al., 2023b), we drew the red bounding box around the interaction element to guide the model’s attention. Additionally, adding the bounding box helped resolve ambiguities at times that the textual action description is not sufficient. For example, sometimes action descriptions are simply empty, and does not contain any description about the element itself. While in other cases, action descriptions exist, but the textual description matches multiple element descriptions and thus results in ambiguity that is only resolved by drawing the bounding box. For example, a button labeled ”Add to Cart” is typically associated with each item in a web shopping list. Without the bounding box, it is impossible to determine which specific button was clicked.
B.4. Android
@@
To address these challenges, we utilized the dataset’s utilities to overlay actions on top of the screenshots, as well as bounding box annotations for post-process detected UI elements. However, we found that the added element annotation marks often confused the model added noise and critical information on the screen. As a result, we proceeded with experiments using a sequence of screenshots that had the actions (tap, drag, and type) drawn over them. AitW’s provided visualization tools to draw the actions, also included labels of special actions such as the back button, home button, and enter. As well as a special ”status” action: either Task Complete or Task Impossible. However, we found that the label Task Complete often confused the models. Despite our efforts to make the models ignore it, we eventually abandoned this specific annotation as it is a technicality of data representation.
@@ Appendix C Error Analysis
Figure 4. An illustration in which the user chose a specific car primarily for its 12-inch feature, but since it was also the cheapest, annotators incorrectly assumed cost was the deciding factor.
[…]
The figure presents an example from the Mind2Web dataset, showcasing four different trucks, each with distinct wheel sizes and prices. The image is annotated with the user’s selection: the second truck, which has the lowest price and features 12” wheels.
@@ C.1. Web
Figure 5. Comparison of ”Match” proportions between Gemini 1.5 Pro and GPT-4-Turbo models across the different domains
[…]
The figure presents a histogram comparing the ”Match” metric values across different domains in the Android in the Zoo dataset. The domains include ”General,” ”Google Apps,” ”Install,” and ”Web Shopping.” The comparison is made between the GPT-4 and Gemini 1.5 Pro models.
C.2. Android
Comparing GPT and Gemini’s performance on the Android dataset revealed notable variations across different domains. Both models were proficient in the “General” domain, but faced challenges in “Web Shopping” and “Install”. This disparity is due to the nature of the “General” dataset, which contains search queries that explicitly reveal the user’s intent. “Google Apps” tasks overlap with PixelHelp, primarily involving settings configuration. Some tasks are ambiguous, such as a button toggle, but the trajectory displays only one option. Additionally, specific tasks request actions that have already been completed, e.g Turn On location history, but the trajectory only shows viewing the Location History setting page, leading to confusion for models. Additionally, the models faced difficulty comprehending the correct order of the sequence of actions that occurred, frequently mistaking the final state (on or off).
@@ C.2. Android
The “Install” domain often presents ambiguous tasks in the format “open (install if not installed)” which confuses both models. Furthermore, in some cases the apps were already pre-installed which made it impossible to predict (model predicts “open”), providing only partial satisfaction in one direction. These results indicate that further refinement and training may be needed to improve the models’ performance in specific domains.
@@ Appendix D Instructions and Prompts
Figure 6. Instructions for human-annotators to conclude if a task is fulfilled by a trajectory, and if two task descriptions satisfying each other. These instructions were the core prompt of the automatic evaluator.
[…]
The figure contains text that was used to instruct human participants on how to evaluate the fulfillment and satisfaction relations when comparing two task descriptions with a UI trajectory.
Figure 7. Goal identification annotation task instruction to measure human performance.
[…]
The figure contains text that was used to instruct human participants while performing the task of identifying user goals from UI trajectories.
Figure 8. Instruction used to guide GPT and Gemini when predicting a task description given a UI trajectory. We swap AitW section and Mind2Web according to the input.
[…]
The figure contains text that was used to guide Gemini 1.5 Pro and GPT-4 models in predicting user goals based on UI trajectories.
Figure 9. Model instructions for evaluating Satisfaction relation between two task descriptions (A and B), given a corresponding trajectory. Includes web/mobile data format instructions and few-shot examples for experimentation (detailed on the next page).
[…]
The figure contains a table with text displaying the instructions used to guide the GPT-4 model in automatically rating the satisfaction relation between two task descriptions.
Figure 10. Few-Shot exemplars demonstrating the Satisfaction Relation Prompt (Figure 9). These simplified examples highlight the nuances of task satisfaction in real-world scenarios.
[…]
The figure contains a table with text, displaying few-shot examples used in the prompt to guide the GPT-4 model in automatically rating two task descriptions.